To the nearest year find the half-life of plutonium -239 given that it decays according to the formula A= A0^e-0.0000284t
Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

A= A0^e-0.0000284t
A=(1/2)*Ao
Therefore 1/2Ao=Ao^e(-0.0000284t)
A0 cancel
take logs of both sides
log (1/2)=-0.0000284t, because ln of e^(t)=t. Ln and e annihilate each other.
-0.693/-0.0000284=24,401 years=t
That compares favorably with quoted 25,000 years.

Answer by MathTherapy(10556)


To the nearest year find the half-life of plutonium -239 given that it decays according to the formula A= A0^e-0.0000284t
t, or time =